### v2

# ReadMe.LLM: A Framework to Help LLMs Understand Your Library

Abstract
Large Language Models (LLMs) often struggle with code generation tasks involving niche software libraries. Existing code generation techniques with only human-oriented documentation can fail —– even when the LLM has access to web search and the library is documented online. To address this challenge, we propose ReadMe.LLM, LLM-oriented documentation for software libraries. By attaching the contents of ReadMe.LLM to a query, performance consistently improves to near-perfect accuracy, with one case study demonstrating up to 100% success across all tested models. We propose a software development lifecycle where LLM-specific documentation is maintained alongside traditional software updates. In this study, we present two practical applications of the ReadMe.LLM idea with diverse software libraries, highlighting that our proposed approach could generalize across programming domains.
1 Introduction
Large Language Models (LLMs) like GPT [1] and Llama [2] have transformed the software development ecosystem. More engineers are using LLMs to generate code with existing software libraries, leveraging these tools to approach coding tasks more efficiently and intuitively. In some cases, we are even seeing AI agents begin to replace human developers themselves. These models, often used as coding assistants, are capable of generating code, debugging, and creating documentation through natural language prompting. One emerging trend, frequently referred to as “vibe coding” [3], involves engineers prompting LLMs with simple, high-level natural language instructions and iteratively refining their code based on the model’s suggestions. This interactive exploratory approach enables fast prototyping and creates a more fluid software development process.
1.1 Challenge
However, not all libraries are equally represented in LLM training data. Well-established libraries like Pandas [4] have plenty of public documentation, Stack Overflow questions, and other resources that are ingested during LLM pretraining, allowing the LLM to produce reliable output, while lesser-known libraries are often misused or misrepresented in AI-generated code [5, 6, 7]. This gap negatively impacts both engineers and library developers. Engineers receive incorrect code, leading to frustration, prolonged debugging, and increased company resource expenditure [8]. Meanwhile, library developers risk losing potential users who abandon their tools in favor of alternatives that work seamlessly with LLM-generated code.
In addition, as AI agents and services become more popular and increasingly integrated into development, their reliance on LLMs amplifies the underrepresentation of smaller libraries. If these agents struggle with less-documented tools, workflows become inefficient, reinforcing a cycle where only well-known libraries thrive.
This dynamic is reshaping the entire software ecosystem. Smaller libraries lose potential users not due to their technical merit but because LLMs fail to capture them accurately. For engineers, this means fewer viable options and slower innovation. Our work addresses these systemic consequences by creating a framework that ensures LLMs can correctly understand and utilize any software library, leveling the playing field and fostering a more accessible development landscape.
1.2 Existing Solutions
There are different techniques for leveraging LLMs for code generation tasks. We list key categories and examples in Figure 1. Many strategies, such as Retrieval Augmented Generation (RAG) [9], require additional infrastructure that falls outside of the typical developer workflow using standard IDEs111Of course, one possible view of the future would involve building a well standardized approach to integrating RAG with IDEs and copilots [10]. But we are not there yet.. The most accessible approach that offers the least friction to development is Prompt Engineering [11]. Within Prompt Engineering, there has been research on thought generation, such as Chain of Thought [12], or how models perform in a “Zero-Shot” [13] manner. We propose a specific prompting framework -– ReadMe.LLM —-which leverages assets (e.g. function signatures, examples, and descriptions) from a respective software library to assist code generation tasks.
Additionally, there are different techniques for delivering updated contexts to an LLM. Recently, continual learning (CL) research has grown to be a good workaround to model cutoff dates. CL enables models to integrate new knowledge without forgetting past information through processes such as multiple training stages [20]. This shift underscores the importance of efficient mechanisms for integrating new knowledge.
In parallel, tools have been developed to automatically extract information from GitHub libraries. Gitingest, a popular tool, automatically extracts the repository directory structure and aggregates its files to be easily copied [21]. This enables users to easily copy entire repositories when trying to prompt LLMs. However, when applying this tool to our case studies, we found that the resulting file was too large and often caused the models to hallucinate.
Libraries aren’t the only tools that can be used as a building block by LLMs. A related approach to providing LLM-specific files is llms.txt, a structured Markdown file organizing website content for LLMs and Agents [22], While web content is primarily designed for human users, this can be restrictive to LLMs with search capabilities or Agents that interact with the web. By providing a concise and structured representation of the content, llms.txt can enhance usability. This takes inspiration from robots.txt files, which detail which URLs a search engine crawler can access. In this context, our ReadMe.LLM proposal extends this idea by offering a well-defined framework for code generation tasks, as opposed to general website content.
1.3 A Novel Elementary Approach: ReadMe.LLM
Current documentation, such as ReadMe.md files, is written for human readers, but LLMs interpret information differently and are less effective with human-targeted formats. We argue that there should be LLM-targeted documentation. To address this, we propose ReadMe.LLM, a structured format to streamline library usage by LLMs:
Optimized Documentation for LLMs: ReadMe.LLM provides structured descriptions of the codebase. Just as traditional header files help tell how to use a library to a traditional compiler, the ReadMe.LLM file tells an LLM how to effectively use this library to get things done.
Seamless Integration: Library developers easily create and attach a ReadMe.LLM to their codebase, which engineers can copy-paste or upload along with their query.
This approach shifts the focus to empowering libraries to be LLM-friendly, fostering adoption of emerging libraries. The overall workflow is illustrated below:
2 ReadMe.LLM
ReadMe.LLM is LLM-oriented documentation: a structured format leveraging assets from a software library to assist code generation tasks. Just as a ReadMe.md file provides essential information to human developers, a ReadMe.LLM provides it to LLMs. Based on our explorative testing and research into prompt engineering strategies, we propose the following ReadMe.LLM structure:
1.
Rules: A customizable set of guidelines that instruct the LLM on how to process the library’s information
2.
Library Description: A concise overview that sets the scene by outlining the library’s purpose, core functionalities, and domain context.
3.
Code snippets: Clear function signatures are paired with illustrative examples that demonstrate real-world usage and expected outcomes.
This was the structure we found worked best based on the libraries we experimented with; however, libraries from different domains may need to make small adjustments. We used XML tags to separate different types of content (e.g. examples). This formatting improves readability for LLMs and helps them easily parse the rules, description, and code snippets in ReadMe.LLM [23].
With ReadMe.LLM defined, we outline how developers can utilize it in the software development process. Below are workflows we envision for three main user personas – library developers, engineers, and AI agents.
2.1 Library Developer
Mirroring the ReadMe.md file for human documentation, we encourage library developers to create a ReadMe.LLM for their libraries to provide LLMs with targeted documentation that enhances coding outcomes. The general process is as follows: a library is released for users on Github, important code snippets and example usage are extracted from that codebase, and this is put together in a formatted text file-–the ReadMe.LLM. Once released, developers can engage with the user community to gather feedback and iterate on the ReadMe.LLM, improving its clarity and effectiveness over time. This falls into the software development cycle. Just as new releases for libraries require updated release notes to inform users of changes, library developers can edit the existing ReadMe.LLM file with any important changes.
2.2 Engineer
Many engineers have turned to an LLM for code generation assistance, but there is no standardized workflow for this process. We define the following general process: the engineers copy the ReadMe.LLM from the library’s repository, they then paste it into an LLM, and finally enter their query. With better context, the LLM provides more accurate and relevant code. If the engineer experiences any pain points, such as a missing function or an unclear example, they can submit this feedback to the library developers.
2.3 AI Agents
AI Agents have become increasingly powerful and represent another user persona that can leverage ReadMe.LLM. Agents are powered by LLMs to automate tasks. Model Context Protocol (MCP) defines a standard way for AI agents to connect to data sources [24, 25]. MCP makes it easier for AI agents to process information and use different information sources when executing a task. Similarly, there has been a rise in the AI Agent libraries and services themselves, which use protocols such as MCP. Browser-Use [26] is a service that enables AI agents to automate tasks within web environments, and Manus [27] is an AI agent that executes tasks autonomously. With AI protocols such as MCP to manage the flow of tasks, Browser-Use and Manus can interact with each other and other tools more efficiently. Agents built using an MCP framework can seamlessly integrate ReadMe.LLM, allowing them to prioritize its contents, maintain context history across different ReadMe.LLM files, and navigate repositories efficiently. With this capability, agents can quickly locate and leverage the relevant ReadMe.LLM when tasked with coding. The process unfolds as follows: the AI Agent identifies a library to use for a task, locates the ReadMe.LLM file and processes it, combines this information with other sources to generate code, and finally, the agent debugs and optimizes the implementation before delivering the final output. This creates a more robust ecosystem where both human engineers and AI agents can utilize diverse libraries with ReadMe.LLM.
3 Experiments
To understand what would be needed in a ReadMe.LLM, we systematically evaluated code that was generated by LLMs using different combinations of software library information. The software library information that we used includes human documentation (ReadMe.md files) and direct code snippets (full-function implementations, usage examples). To ensure robustness, we tested this across five different LLMs, all accessed through Perplexity: GPT-4o, Sonar Huge (built on top of LLaMA 3.3 70B), Claude 3.7 Sonnet, Grok-2, and Deepseek R1. However, during our experimentation, DeepSeek R1 was temporarily removed from Perplexity, so we completed its testing via the DeepSeek website.
Something to consider is that LLMs have a knowledge cutoff, meaning they lack awareness of new information beyond their last training date (Table 1), and high training costs prevent frequent updates [28]. Since large-scale continual learning is still an open challenge, we relied on web search–which aggregates information from a broad range of sources [29]—as a practical alternative for accessing up-to-date information. This reflects realistic scenarios where users seek the most current insights. Additionally, to evaluate ReadMe.LLM’s utility in settings where web search is not feasible, such as internal company libraries, we included iterations without web search.
| Model | Cutoff Date |
| GPT-4o | October 2023 [30, 31] |
| Llama 3.3 70B | December 2023 [32, 31] |
| Claude 3.7 Sonnet | April 2024 [33, 31] |
| Grok-2 | July 2024 [34, 35] |
| Deepseek R1 | July 2024 [36, 37] |
With these LLMs, we experimented with two distinct libraries: DigitalRF [38] and Supervision [39]. DigitalRF, an academic library with limited documentation, represents libraries that are not likely to have been included in the LLMs’ training process. Supervision, a modern, industry-run library, helps assess whether similar limitations persist for newer but more widely used libraries.
For each library, we designed tasks based on consultations with the library developers to get insight into common use cases, ensuring LLMs interact with them realistically. We then provided these tasks to the LLMs, collected their generated code, and evaluated performance using two criteria:
1.
Minimal Debugging – Code should work with at most three debugging rounds; the user pastes the error and the LLM regenerates fixed code based on that.
2.
Correct Library Utilization – The LLM should use the intended library functions rather than recreating functionality from scratch.
After evaluating which context combinations yielded the highest success, we developed an optimal ReadMe.LLM for each library that generalizes well. We verified our optimal ReadMe.LLM on a held-out test using two previously untested LLMs –Gemini 2.0 Flash and Mistral Large (Cutoff dates: Dec 2024 [31] and 2023 [40], respectively).
3.1 Finding the Optimal ReadMe.LLM
3.1.1 Case Study 1: Supervision
Supervision [39] is an industry-led library developed by Roboflow, which simplifies the process of working with computer vision models. It offers connectors to popular model libraries, a plethora of visualizers (annotators), powerful post processing features, and an easy learning curve. Main capabilities include: Detect and Annotate, Save Detections, Filter Detections, Detect Small Objects, Track Objects on Video, and Process Datasets.
Experiment Process
For Supervision, we tasked LLMs with detecting, annotating, and cropping cars in an image. We selected an image with multiple objects (such as people, cars and buildings) to introduce complexity and tested the LLMs’ ability to generate code that differentiates between relevant and irrelevant detections.
The LLM had to identify all cars, add a confidence score annotation, save the bounding box coordinates, and crop each detected car. To meet the Correct Library Utilization we mentioned above, the LLM should use Supervision’s Detections, Annotators, and Image Utility classes and functions, rather than alternative libraries and methods. The code should return at least one cropped image of a car, and the annotated picture should show confidence through either a bar or a percentage.
Results
Figure 7 highlights that adding any context significantly improves LLM performance. The baseline success rate without context averaged around 30%. Interestingly, DeepSeek R1 saw a decrease in performance when only given ReadMe.md as context – a potential sign that LLMs do not respond well to human-facing documentation. Relying solely on examples achieved a 96% average success rate, while incorporating combined contexts enabled all models to hit 100%, with the exception of Grok-2, which performed notably worse.
3.1.2 Case Study 2: DigitalRF
DigitalRF [38] is an academic library developed by MIT Haystack that encompasses a standardized HDF5 format for reading and writing radio frequency (RF) data. Main capabilities include writing (converting an input WAV file into HDF5 format), and reading (converting HDF5 format back into a WAV file).
DigitalRF presents an interesting contrast from Supervision. It is less popular and has minimal documentation. It is an example of a common class of libraries focused on file format translation, so testing it can help us assess the ReadMe.LLM idea in a broader context.
Experiment Process
For DigitalRF, we tasked the LLMs with writing a WAV file into DigitalRF-HDF5 format. We obtained a WAV file (a 10-second-long radio signal) containing I/Q data using a Software Defined Radio (SDR) and the SDR++ application, and tasked LLMs with converting it to a standardized HDF5 format using the DigitalRF library. To meet the correct library utilization requirement above, we made sure the LLM-generated code created a proper HDF5 folder structure. We ran this output through a pre-built script to reconstruct the original WAV file and ensured that it played back the original audio sample.
Results
Similar to Supervision, we again see in Figure 9 that adding any context significantly improves LLM performance when generating code for unfamiliar libraries. The poor performance with ReadMe.md further proves that LLMs do not respond well to documentation that is intentionally made to be readable to humans.
Incorporating structured information consistently led to better results. Among individual contexts, function-related information and examples had the strongest impact, both raising the average success rate to 64%. For combined contexts, ReadMe.md + Functions achieved the highest success rate at an average of 88%.
3.2 Verifying the ReadMe.LLM
After analyzing the case study results presented earlier, we initiated the design of an optimal ReadMe.LLM for both libraries. The Supervision case study results revealed that using solely the ReadMe.md context led to lower accuracy than when no context was provided. Consequently, we decided to omit ReadMe.md information from our final ReadMe.LLM and instead included only code snippets—interweaving function implementations and code examples.
In the first version of Supervision’s ReadMe.LLM, we incorporated the complete Detections class, all Annotator classes, Image Utility functions, and corresponding examples. We tested this version against Sonar and Grok-2, the models that had previously underperformed. After several iterations, it became evident that this initial version’s extensive length led to hallucinations.
To reduce the length of the context, we revised the ReadMe.LLM by including examples and only function signatures, rather than full implementations. This final version achieved a 100% success rate with Sonar and Grok-2; subsequent testing with GPT-4o, Claude 3.7 Sonnet, and DeepSeek R1 also yielded perfect performance when web search was enabled. When evaluated without web search, the ReadMe.LLM maintained this performance across all models, with the exception of Grok-2, which achieved an 80% success rate.
To verify ReadMe.LLM generalizes to other Supervision use cases, we designed a second task. In this task, the LLM was required to identify individuals within an image, apply a blur to each person, and overlay a different image on each subject.
As shown in Figure 12, all LLMs performed poorly in zero-shot coding—even with web search activated—with only DeepSeek R1 occasionally succeeding. However, when the ReadMe.LLM was provided, the success rate increased to 100% across all models, demonstrating its adaptability to a variety of tasks.
Subsequently, we employed an analogous approach to construct DigitalRF’s ReadMe.LLM, directly interweaving function signatures and examples from the repository, similar to our process for Supervision. This approach immediately yielded a 100% success rate for Sonar and Grok-2 when web search was enabled, and was therefore adopted as our final ReadMe.LLM for DigitalRF. Among the remaining three models, only GPT-4o did not achieve perfect performance with web search, attaining only 80%. When web search was disabled, the average success rate dropped to 70%, with only DeepSeek R1 maintaining a 100% success rate. These results suggest that further refinements could yield an even more effective ReadMe.LLM for DigitalRF in future iterations.
3.3 Held Out Tests
To further assess the robustness of ReadMe.LLM, we conducted a held-out test using two new models that had not been used in previous experiments: Gemini 2.0 Flash and Mistral Large. This evaluation aimed to verify the effectiveness of our final framework when applied to LLMs that did not contribute to our development process.
We began with Supervision. Even with web search enabled, zero-context prompting performed poorly. Gemini succeeded only once out of five trials on the first task, and both models failed entirely on the second. With ReadMe.LLM Gemini’s success rate jumped to 100% on the first task and 80% on the second. Mistral, which had previously failed both, reached a perfect 100% on both tasks. Without web search, Supervision’s ReadMe.LLM sustained a high success rate with both models, with only Gemini exhibiting a slight decline.
We tested these same models with DigitalRF. When prompted without any additional context, accuracy was 0%, even with web search capabilities. Once ReadMe.LLM was applied, however, both models achieved an 80% success rate, showing a dramatic and consistent improvement. In line with the original models, the absence of web search capabilities resulted in a slight performance decline, but still remained significantly superior to conditions without context.
This consistent performance boost demonstrates that ReadMe.LLM significantly enhances code generation capabilities, even for held-out models and libraries that did not influence the development of ReadMe.LLM. Our results affirm that ReadMe.LLM not only improves accuracy with familiar models, but also generalizes effectively to entirely new ones. By bridging gaps in LLM knowledge, ReadMe.LLM makes code generation more reliable and robust across diverse architectures and domains.
4 Discussion
It’s Possible to Seamlessly Improve Code Generation Through Prompting with LLM-Oriented Documentation. Through our experimentation and research, we show that prompting with LLM-oriented software library documentation —ReadMe.LLM– can greatly increase the performance of code completion tasks for LLMs. ReadMe.LLM can perform optimally in the majority of scenarios irrespective of model selection. This is a powerful tool for library developers and engineers, because it increases the accessibility and performance of leveraging LLMs for code completion. Engineers no longer have to create complex prompts, use computationally-intensive methods like RAG, or drastically change their queries. Instead, they can just attach a library’s ReadMe.LLM and ask questions as usual. Similarly, library developers can develop a ReadMe.LLM to ensure that their library is being correctly represented by an LLM and therefore, seamlessly used by the targeted engineer.
Tailoring context selection to the model can improve code quality. Through our experimentation, we have found that different models have varying success with diverse contexts. While we have identified a framework that consistently performs well across tested models, there may be situations where you can gain even better performance by modifying the library’s ReadMe.LLM. For example, in our DigitalRF case study, we observed that Sonar achieved a 100% success rate with ReadMe.md and Examples, but dropped to 80% with ReadMe.md and Functions. However, the other models (Grok-2, GPT-4o, Claude 3.7, and DeepSeek R1) saw better performance with ReadMe.md + Functions. With this in mind, it may be advisable for library developers to test their ReadMe.LLM against a wide variety of LLMs to ensure its robustness.
Patterns in Models’ Limitations. There are several challenges an LLM would face when completing our tasks. After our experimentation, we were able to categorize these challenges into LLM code generation insights.
First, it became a common occurrence that a model would fail on a task, not because of the usage of the target library, but because of the prerequisite Input/Output tasks. Errors such as importing a suitable library, creating a new file to save data, or reading the correct data often caused the task to fail. For example, with DigitalRF, the model failed at reading in the WAV file because the LLM-generated code utilized the wrong Python library that supported a different file format. We argue that this is a general reflection of an LLM’s ability to generate code for IO-related tasks. With this, if a library wants to enhance developer use, developers should provide necessary context about IO-tasks in the ReadMe.LLM.
Second, the model would often hallucinate by using an alternative method or library, failing to use the targeted library’s code. For example, with Supervision, the model would often crop images by slicing a dictionary, rather than using the crop image function defined in Supervision. This highlights the importance of adding examples and function definitions, especially when the functionality may conflict with existing libraries that a model is trained on. This can help ground the model in using the correct function from the targeted library.
5 Conclusion and Future Work
We present ReadMe.LLM, novel LLM-oriented documentation that provides relevant context about a software library to assist code generation. We evaluated different combinations and structures of context and tested these across the current leading LLMs. We presented the optimal ReadMe.LLM structure, which has the highest average accuracy across different models, and increases correctness by 5x.
As engineers continue to turn to LLMs when facing roadblocks, library developers must make their content easily available and understandable to LLMs. Failure to do so will not only hinder engineers by producing unreliable code but also disadvantage smaller libraries, perpetuating a cycle of underutilization and inefficiency. ReadMe.LLM becomes essential for bridging the gap between library documentation and Generative AI assistance.
Looking ahead, we remain committed to enhancing this framework by exploring new components and optimizing the structure of context delivery. We are planning on investigating how this framework extends to tasks other than code generation, such as question & answering and code debugging. Additionally, in this paper, we focus on LLM chatbots, but ReadMe.LLM can be extended to co-pilots as well. With the rise of vibe-coding and the adoption of products like Cursor [41], improving the code generation capabilities directly in the editor is important. A co-pilot could recognize the ReadMe.LLM file within an imported module’s directory and utilize it to generate more relevant and accurate code for its user.
We welcome contributions from the community to advance this initiative and shape the future of LLM-library interactions. Please explore our website, readmellm.github.io, and we encourage you to contribute to online discussions.
Acknowledgments
We thank the National Science Foundation and especially the SpectrumX (AST-2132700) community for its support. We also thank the UC Berkeley College of Engineering’s Fung Institute, as well as Dr. Josh Sanz for helpful conversations.