### v3

# ReadMe.LLM: A Framework to Help LLMs Understand Your Library

Abstract
Large Language Models (LLMs) often struggle with code generation tasks involving niche software libraries. Existing code generation techniques with only human-oriented documentation can fail – even when the LLM has access to web search and the library is documented online. To address this challenge, we propose ReadMe.LLM, LLM-oriented documentation for software libraries. By attaching the contents of ReadMe.LLM to a query, performance consistently improves to near-perfect accuracy, with one case study demonstrating up to 100% success across all tested models. We propose a software development lifecycle where LLM-specific documentation is maintained alongside traditional software updates. In this study, we present two practical applications of the ReadMe.LLM idea with diverse software libraries, highlighting that our proposed approach could generalize across programming domains.
1 Introduction
Large Language Models (LLMs) like GPT [1] and Llama [2] have transformed the software development ecosystem. More engineers are using LLMs to generate code with existing software libraries, leveraging these tools to approach coding tasks more efficiently and intuitively. In some cases, we are even seeing AI agents begin to replace human developers themselves. These models, often used as coding assistants, are capable of generating code, debugging, and creating documentation through natural language prompting.
These advances build on a growing lineage of code-specific language models. OpenAI’s Codex [3], fine-tuned on GitHub code, enabled natural language-to-code translation and powers tools like GitHub Copilot [4]. Meta’s Code LLaMA [5] was pretrained specifically on code-related data, allowing it to support multiple programming languages and longer context windows. Both models expanded the capabilities of LLMs while maintaining low-friction interfaces for developers — but they also highlight a core challenge: for this innovation and performance to continue, LLMs rely on consistent and structured documentation. Recent work has emphasized that structured specification (precise descriptions of a component’s expected behavior, inputs, and outputs) is essential to making LLM-based systems more modular and reliable [6]. As these systems become increasingly integrated into IDEs and developer workflows, new coding practices continue to emerge.
One emerging trend, frequently referred to as “vibe coding” [7], involves engineers prompting LLMs with simple, high-level natural language instructions and iteratively refining their code based on the model’s suggestions. This interactive exploratory approach enables fast prototyping and creates a more fluid software development process.
1.1 Challenge
However, not all libraries are equally represented in LLM training data. Well-established libraries like Pandas [8] have plenty of public documentation, Stack Overflow questions, and other resources that are ingested during LLM pretraining, allowing the LLM to produce reliable output, while lesser-known libraries are often misused or misrepresented in AI-generated code [9, 10, 11]. This gap negatively impacts both engineers and library developers. Engineers receive incorrect code, leading to frustration, prolonged debugging, and increased company resource expenditure [12]. Meanwhile, library developers risk losing potential users who abandon their tools in favor of alternatives that work seamlessly with LLM-generated code.
In addition, as AI agents and services become more popular and increasingly integrated into development, their reliance on LLMs amplifies the underrepresentation of smaller libraries. If these agents struggle with less-documented tools, workflows become inefficient, reinforcing a cycle where only well-known libraries thrive.
This dynamic is reshaping the entire software ecosystem. Smaller libraries lose potential users not due to their technical merit but because LLMs fail to capture them accurately. For engineers, this means fewer viable options and slower innovation. Our work addresses these systemic consequences by creating a framework that ensures LLMs can correctly understand and utilize any software library, leveling the playing field and fostering a more accessible development landscape.
1.2 Existing Solutions
There are different techniques for leveraging LLMs for code generation tasks. We list key categories and examples in Figure 1. Many strategies, such as Retrieval Augmented Generation (RAG) [13], require additional infrastructure that falls outside of the typical developer workflow using standard IDEs111Of course, one possible view of the future would involve building a well standardized approach to integrating RAG with IDEs and copilots [14]. But we are not there yet.. The most accessible approach that offers the least friction to development is Prompt Engineering [15]. Within Prompt Engineering, there has been research on thought generation, such as Chain of Thought [16], or how models perform in a “Zero-Shot” [17] manner. CodeScholar [24] addresses this by generating realistic, idiomatic API usage examples to help developers understand how to use unfamiliar libraries, serving as a complement to traditional documentation. These examples not only improve human comprehension but also improve the quality of LLM-generated code when used in downstream tasks. We propose a specific prompting framework -– ReadMe.LLM —-which leverages assets (e.g. function signatures, examples, and descriptions) from a respective software library to assist code generation tasks.
Additionally, there are different techniques for delivering updated contexts to an LLM. Recently, continual learning (CL) research has grown to be a good workaround to model cutoff dates. CL enables models to integrate new knowledge without forgetting past information through processes such as multiple training stages [25]. This shift underscores the importance of efficient mechanisms for integrating new knowledge.
In parallel, tools have been developed to automatically extract information from GitHub libraries. Gitingest, a popular tool, automatically extracts the repository directory structure and aggregates its files to be easily copied [26]. This enables users to easily copy entire repositories when trying to prompt LLMs. However, when applying this tool to our case studies, we found that the resulting file was too large and often caused the models to hallucinate.
Libraries aren’t the only tools that can be used as a building block by LLMs. A related approach to providing LLM-specific files is llms.txt, a structured Markdown file organizing website content for LLMs and Agents [27], While web content is primarily designed for human users, this can be restrictive to LLMs with search capabilities or Agents that interact with the web. By providing a concise and structured representation of the content, llms.txt can enhance usability. This takes inspiration from robots.txt files, which detail which URLs a search engine crawler can access. In this context, our ReadMe.LLM proposal extends this idea by offering a well-defined framework for code generation tasks, as opposed to general website content.
1.3 A Novel Elementary Approach: ReadMe.LLM
Current documentation, such as ReadMe.md files, is written for human readers, but LLMs interpret information differently and are less effective with human-targeted formats. We argue that there should be LLM-targeted documentation 222As we were writing this report after completing our experiments, Andrej Karpathy tweeted: ”It’s 2025 and most content is still written for humans instead of LLMs. 99.9% of attention is about to be LLM attention, not human attention. E.g. 99% of libraries still have docs that basically render to some pretty .html static pages assuming a human will click through them. In 2025 the docs should be a single your_project.md text file that is intended to go into the context window of an LLM. Repeat for everything.”[28].. To address this, we propose ReadMe.LLM, a structured format to streamline library usage by LLMs:
Optimized Documentation for LLMs: ReadMe.LLM provides structured descriptions of the codebase. Just as traditional header files help tell how to use a library to a traditional compiler, the ReadMe.LLM file tells an LLM how to effectively use this library to get things done.
Seamless Integration: Library developers easily create and attach a ReadMe.LLM to their codebase, which engineers can copy-paste or upload along with their query.
This approach shifts the focus to empowering libraries to be LLM-friendly, fostering adoption of emerging libraries. The overall workflow is illustrated below:
2 ReadMe.LLM
ReadMe.LLM is LLM-oriented documentation: a structured format leveraging assets from a software library to assist code generation tasks. Just as a ReadMe.md file provides essential information to human developers, a ReadMe.LLM provides it to LLMs. Based on our explorative testing and research into prompt engineering strategies, we propose the following ReadMe.LLM structure:
1.
Rules: A customizable set of guidelines that instruct the LLM on how to process the library’s information
2.
Library Description: A concise overview that sets the scene by outlining the library’s purpose, core functionalities, and domain context.
3.
Code snippets: Clear function signatures are paired with illustrative examples that demonstrate real-world usage and expected outcomes.
This was the structure we found worked best based on the libraries we experimented with; however, libraries from different domains may need to make small adjustments. We used XML tags to separate different types of content (e.g. examples). This formatting improves readability for LLMs and helps them easily parse the rules, description, and code snippets in ReadMe.LLM [29]. The complete ReadMe.LLM docs used in our experiments can be found in Appendix C. In order to create a ReadMe.LLM for a library follow the instructions in Appendix B.
With ReadMe.LLM defined, we outline how developers can utilize it in the software development process. Below are workflows we envision for three main user personas – library developers, engineers, and AI agents.
2.1 Library Developer
Mirroring the ReadMe.md file for human documentation, we encourage library developers to create a ReadMe.LLM for their libraries to provide LLMs with targeted documentation that enhances coding outcomes. The general process is as follows: a library is released for users on Github, important code snippets and example usage are extracted from that codebase, and this is put together in a formatted text file-–the ReadMe.LLM. Once released, developers can engage with the user community to gather feedback and iterate on the ReadMe.LLM, improving its clarity and effectiveness over time. This falls into the software development cycle. Just as new releases for libraries require updated release notes to inform users of changes, library developers can edit the existing ReadMe.LLM file with any important changes.
2.2 Engineer
Many engineers have turned to an LLM for code generation assistance, but there is no standardized workflow for this process. We define the following general process: the engineers copy the ReadMe.LLM from the library’s repository, they then paste it into an LLM, and finally enter their query. With better context, the LLM provides more accurate and relevant code. If the engineer experiences any pain points, such as a missing function or an unclear example, they can submit this feedback to the library developers.
2.3 AI Agents
AI Agents have become increasingly powerful and represent another user persona that can leverage ReadMe.LLM. Agents are powered by LLMs to automate tasks. Model Context Protocol (MCP) defines a standard way for AI agents to connect to data sources [30, 31]. MCP makes it easier for AI agents to process information and use different information sources when executing a task. Similarly, there has been a rise in the AI Agent libraries and services themselves, which use protocols such as MCP. Browser-Use [32] is a service that enables AI agents to automate tasks within web environments, and Manus [33] is an AI agent that executes tasks autonomously. With AI protocols such as MCP to manage the flow of tasks, Browser-Use and Manus can interact with each other and other tools more efficiently. Agents built using an MCP framework can seamlessly integrate ReadMe.LLM, allowing them to prioritize its contents, maintain context history across different ReadMe.LLM files, and navigate repositories efficiently. With this capability, agents can quickly locate and leverage the relevant ReadMe.LLM when tasked with coding. The process unfolds as follows: the AI Agent identifies a library to use for a task, locates the ReadMe.LLM file and processes it, combines this information with other sources to generate code, and finally, the agent debugs and optimizes the implementation before delivering the final output. This creates a more robust ecosystem where both human engineers and AI agents can utilize diverse libraries with ReadMe.LLM.
3 Experiments
To understand what would be needed in a ReadMe.LLM, we systematically evaluated code that was generated by LLMs using different combinations of software library information. The software library information that we used includes human documentation (ReadMe.md files) and direct code snippets (full-function implementations, usage examples). To ensure robustness, we tested this across five different LLMs, all accessed through Perplexity: GPT-4o, Sonar Huge (built on top of LLaMA 3.3 70B), Claude 3.7 Sonnet, Grok-2, and Deepseek R1. However, during our experimentation, DeepSeek R1 was temporarily removed from Perplexity, so we completed its testing via the DeepSeek website.
Something to consider is that LLMs have a knowledge cutoff, meaning they lack awareness of new information beyond their last training date (Table 1), and high training costs prevent frequent updates [34]. Since large-scale continual learning is still an open challenge, we relied on web search–which aggregates information from a broad range of sources [35]—as a practical alternative for accessing up-to-date information. This reflects realistic scenarios where users seek the most current insights. Additionally, to evaluate ReadMe.LLM’s utility in settings where web search is not feasible, such as internal company libraries, we included iterations without web search.
| Model | Cutoff Date |
| GPT-4o | October 2023 [36, 37] |
| Llama 3.3 70B | December 2023 [38, 37] |
| Claude 3.7 Sonnet | April 2024 [39, 37] |
| Grok-2 | July 2024 [40, 41] |
| Deepseek R1 | July 2024 [42, 43] |
With these LLMs, we experimented with two distinct libraries: DigitalRF [44] and Supervision [45]. DigitalRF, an academic library with limited documentation, represents libraries that are not likely to have been included in the LLMs’ training process. Supervision, a modern, industry-run library, helps assess whether similar limitations persist for newer but more widely used libraries.
For each library, we designed tasks based on consultations with the library developers to get insight into common use cases, ensuring LLMs interact with them realistically. We then provided these tasks to the LLMs, collected their generated code, and evaluated performance using two criteria:
1.
Minimal Debugging – Code should work with at most three debugging rounds; the user pastes the error and the LLM regenerates fixed code based on that.
2.
Correct Library Utilization – The LLM should use the intended library functions rather than recreating functionality from scratch.
After evaluating which context combinations yielded the highest success, we developed an optimal ReadMe.LLM for each library that generalizes well. We verified our optimal ReadMe.LLM on a held-out test using two previously untested LLMs –Gemini 2.0 Flash and Mistral Large (Cutoff dates: Dec 2024 [37] and 2023 [46], respectively).
3.1 Finding the Optimal ReadMe.LLM
3.1.1 Case Study 1: Supervision
Supervision [45] is an industry-led library developed by Roboflow, which simplifies the process of working with computer vision models. It offers connectors to popular model libraries, a plethora of visualizers (annotators), powerful post processing features, and an easy learning curve. Main capabilities include: Detect and Annotate, Save Detections, Filter Detections, Detect Small Objects, Track Objects on Video, and Process Datasets.
Experiment Process
For Supervision, we tasked LLMs with detecting, annotating, and cropping cars in an image. We selected an image with multiple objects (such as people, cars and buildings) to introduce complexity and tested the LLMs’ ability to generate code that differentiates between relevant and irrelevant detections.
The LLM had to identify all cars, add a confidence score annotation, save the bounding box coordinates, and crop each detected car. To meet the Correct Library Utilization we mentioned above, the LLM should use Supervision’s Detections, Annotators, and Image Utility classes and functions, rather than alternative libraries and methods. The code should return at least one cropped image of a car, and the annotated picture should show confidence through either a bar or a percentage.
Results
Figure 7 highlights that adding any context significantly improves LLM performance. The baseline success rate without context averaged around 30%. Interestingly, DeepSeek R1 saw a decrease in performance when only given ReadMe.md as context – a potential sign that LLMs do not respond well to human-facing documentation. Relying solely on examples achieved a 96% average success rate, while incorporating combined contexts enabled all models to hit 100%, with the exception of Grok-2, which performed notably worse.
3.1.2 Case Study 2: DigitalRF
DigitalRF [44] is an academic library developed by MIT Haystack that encompasses a standardized HDF5 format for reading and writing radio frequency (RF) data. Main capabilities include writing (converting an input WAV file into HDF5 format), and reading (converting HDF5 format back into a WAV file).
DigitalRF presents an interesting contrast from Supervision. It is less popular and has minimal documentation. It is an example of a common class of libraries focused on file format translation, so testing it can help us assess the ReadMe.LLM idea in a broader context.
Experiment Process
For DigitalRF, we tasked the LLMs with writing a WAV file into DigitalRF-HDF5 format. We obtained a WAV file (a 10-second-long radio signal) containing I/Q data using a Software Defined Radio (SDR) and the SDR++ application, and tasked LLMs with converting it to a standardized HDF5 format using the DigitalRF library. To meet the correct library utilization requirement above, we made sure the LLM-generated code created a proper HDF5 folder structure. We ran this output through a pre-built script to reconstruct the original WAV file and ensured that it played back the original audio sample.
Results
Similar to Supervision, we again see in Figure 9 that adding any context significantly improves LLM performance when generating code for unfamiliar libraries. The poor performance with ReadMe.md further proves that LLMs do not respond well to documentation that is intentionally made to be readable to humans.
Incorporating structured information consistently led to better results. Among individual contexts, function-related information and examples had the strongest impact, both raising the average success rate to 64%. For combined contexts, ReadMe.md + Functions achieved the highest success rate at an average of 88%.
3.2 Verifying the ReadMe.LLM
After analyzing the case study results presented earlier, we initiated the design of an optimal ReadMe.LLM for both libraries. The Supervision case study results revealed that using solely the ReadMe.md context led to lower accuracy than when no context was provided. Consequently, we decided to omit ReadMe.md information from our final ReadMe.LLM and instead included only code snippets—interweaving function implementations and code examples.
In the first version of Supervision’s ReadMe.LLM, we incorporated the complete Detections class, all Annotator classes, Image Utility functions, and corresponding examples. We tested this version against Sonar and Grok-2, the models that had previously underperformed. After several iterations, it became evident that this initial version’s extensive length led to hallucinations.
To reduce the length of the context, we revised the ReadMe.LLM by including examples and only function signatures, rather than full implementations. This final version achieved a 100% success rate with Sonar and Grok-2; subsequent testing with GPT-4o, Claude 3.7 Sonnet, and DeepSeek R1 also yielded perfect performance when web search was enabled. When evaluated without web search, the ReadMe.LLM maintained this performance across all models, with the exception of Grok-2, which achieved an 80% success rate.
To verify ReadMe.LLM generalizes to other Supervision use cases, we designed a second task. In this task, the LLM was required to identify individuals within an image, apply a blur to each person, and overlay a different image on each subject.
As shown in Figure 12, all LLMs performed poorly in zero-shot coding—even with web search activated—with only DeepSeek R1 occasionally succeeding. However, when the ReadMe.LLM was provided, the success rate increased to 100% across all models, demonstrating its adaptability to a variety of tasks.
Subsequently, we employed an analogous approach to construct DigitalRF’s ReadMe.LLM, directly interweaving function signatures and examples from the repository, similar to our process for Supervision. This approach immediately yielded a 100% success rate for Sonar and Grok-2 when web search was enabled, and was therefore adopted as our final ReadMe.LLM for DigitalRF. Among the remaining three models, only GPT-4o did not achieve perfect performance with web search, attaining only 80%. When web search was disabled, the average success rate dropped to 70%, with only DeepSeek R1 maintaining a 100% success rate. These results suggest that further refinements could yield an even more effective ReadMe.LLM for DigitalRF in future iterations.
A comparison of results for earlier model versions across all contexts is provided in Appendix A
3.3 Held Out Tests
To further assess the robustness of ReadMe.LLM, we conducted a held-out test using three new models that had not been used in previous experiments: Gemini 2.0 Flash, Mistral Large, and GPT-4.1. This evaluation aimed to verify the effectiveness of our final framework when applied to LLMs that did not contribute to our development process.
We began with Supervision. Even with web search enabled, zero-context prompting performed poorly. With ReadMe.LLM Gemini’s success rate jumped to 100% on the first task and 80% on the second. Mistral, which had previously failed both, reached a perfect 100% on both tasks, as did GPT-4.1. Without web search, Supervision’s ReadMe.LLM sustained a high success rate with these models, with only Gemini exhibiting a slight decline.
We tested these same models with DigitalRF. When prompted without any additional context, accuracy was 0%, even with web search capabilities. However, once ReadMe.LLM was applied, Gemini and Mistral achieved a 80% success rate and GPT-4.1 100%, showing a dramatic and consistent improvement. In line with the original models, the absence of web search capabilities resulted in a slight performance decline with Gemini and Mistral, but still remained significantly superior to conditions without context.
This consistent performance boost demonstrates that ReadMe.LLM significantly enhances code generation capabilities, even for held-out models and libraries that did not influence the development of ReadMe.LLM. Our results affirm that ReadMe.LLM not only improves accuracy with familiar models, but also generalizes effectively to entirely new ones. By bridging gaps in LLM knowledge, ReadMe.LLM makes code generation more reliable and robust across diverse architectures and domains.
4 Discussion
It’s Possible to Seamlessly Improve Code Generation Through Prompting with LLM-Oriented Documentation. Through our experimentation and research, we show that prompting with LLM-oriented software library documentation —ReadMe.LLM– can greatly increase the performance of code completion tasks for LLMs. ReadMe.LLM can perform optimally in the majority of scenarios irrespective of model selection. This is a powerful tool for library developers and engineers, because it increases the accessibility and performance of leveraging LLMs for code completion. Engineers no longer have to create complex prompts, use computationally-intensive methods like RAG, or drastically change their queries. Instead, they can just copy and paste a library’s ReadMe.LLM into an LLM’s prompt window and give a code generation task as usual 333After our paper was first released, Andrej Karpathy tweeted in support of the seamless experience of copy and pasting documentation for LLMs: ”Tired: elaborate docs pages for your product/service/library with fancy color palettes, branding, animations, transitions, dark mode, … Wired: one single docs .md file and a ’copy to clipboard’ button” [47].. Similarly, library developers can develop a ReadMe.LLM to ensure that their library is being correctly represented by an LLM and therefore, seamlessly used by the targeted engineer.
Tailoring context selection to the model can improve code quality. Through our experimentation, we have found that different models have varying success with diverse contexts. While we have identified a framework that consistently performs well across tested models, there may be situations where you can gain even better performance by modifying the library’s ReadMe.LLM. For example, in our DigitalRF case study, we observed that Sonar achieved a 100% success rate with ReadMe.md and Examples, but dropped to 80% with ReadMe.md and Functions. However, the other models (Grok-2, GPT-4o, Claude 3.7, and DeepSeek R1) saw better performance with ReadMe.md + Functions. With this in mind, it may be advisable for library developers to test their ReadMe.LLM against a wide variety of LLMs to ensure its robustness.
Patterns in Models’ Limitations. There are several challenges an LLM would face when completing our tasks. After our experimentation, we were able to categorize these challenges into LLM code generation insights.
First, it became a common occurrence that a model would fail on a task, not because of the usage of the target library, but because of the prerequisite Input/Output tasks. Errors such as importing a suitable library, creating a new file to save data, or reading the correct data often caused the task to fail. For example, with DigitalRF, the model failed at reading in the WAV file because the LLM-generated code utilized the wrong Python library that supported a different file format. We argue that this is a general reflection of an LLM’s ability to generate code for IO-related tasks. With this, if a library wants to enhance developer use, developers should provide necessary context about IO-tasks in the ReadMe.LLM.
Second, some models are able to complete the task but do not use the functions provided by the specified library. We marked these occurrences as failures, since utilizing the actual library functions is typically much more efficient than assembling a solution from a mix of other libraries. For example, during our Supervision tasks, the library offered functions to crop or overlay an image, but the LLM generated code that performed these tasks using the CV library instead, resulting in much longer and less efficient code. This is also important from a Software Bill of Materials (SBOM) [48] perspective: an SBOM is an inventory of all software components and dependencies, and using the correct library functions ensures that the SBOM accurately reflects the software’s dependencies, improving transparency and security. This highlights the importance of providing clear examples and function definitions, especially when the desired functionality may overlap with existing libraries that a model is trained on. Doing so can help guide the model to use the correct function from the targeted library.
LLM Performance is a Moving Target. While our final results demonstrate consistent gains using ReadMe.LLM, earlier versions of the same models – shown in Appendix A – showed far more dramatic improvements.
This raises a broader question: as models evolve, which challenges are temporary and which are more fundamental? For example, hallucinations might fade as models improve, or they might persist in new forms. Recent work from Nexusflow on Athene-V2 suggests that models often face a tradeoff: being fine-tuned for conversational chat vs. agentic behavior [49]. This specialization means models may not be able to excel at all tasks simultaneously. And regardless of progress, smaller local models will likely continue to face more limitations due to resource constraints. All of this reinforces the idea that LLM evaluation must keep pace with a fast-moving target.
Jailbreaking and Prompt Robustness. Recent safety research underscores that LLM behavior can be dramatically altered through structured prompting alone. For instance, HiddenLayer’s KROP technique bypasses safety safeguards simply by reformatting the input using specific structural patterns, without needing to inject unsafe content [50]. Inspired by this behavior, ReadMe.LLM encloses library information in XML tags – a way that LLMs are highly responsive to. However, this strength may also be a weakness: such formatting has recently been characterized as a form of ”jailbreaking,” since many models appear overly obedient to syntactically clean inputs like XML. If future alignment interventions are trained to block these structured jailbreaks, ReadMe.LLM’s effectiveness could degrade – even if its content remains entirely safe and constructive. This highlights the need to monitor prompt performance over time and investigate how safety tuning may unintentionally interfere with developer-facing tools.
Similar Approaches in Practical Evaluation. The Berkeley Function-Calling Leaderboard has implemented data filtering and quality enhancement techniques when using live, user-contributed function documentation and queries to benchmark LLM performance in function calling [51, 52]. Their definition of high-quality documentation—structured JSON containing key fields such as function name, description, and detailed specifications—closely aligns with the design of ReadMe.LLM, proving the generalizability of our structure. Furthermore, their evaluation includes comprehensive benchmarking across programming languages and libraries, incorporating methods such as function relevance detection and Abstract Syntax Tree (AST) analysis. Integrating these evaluation techniques into the ReadMe.LLM assessment could enhance its robustness to benchmark and highlight its practical benefits.
In effect, the new thing we are advocating is that every library developer should learn from the benchmark design, just as LLM developers learn from benchmarks. Knowing that LLM developers are chasing higher benchmark scores also informs library developers. While perhaps large popular libraries have the luxury of expecting LLM-providers to optimize for them, smaller niche libraries would do well to make themselves look like the benchmarks that LLM providers optimize for.
Extending ReadMe.LLM to Tool Use. Recent work has increasingly focused on LLMs’ ability to interact with external tools, such as invoking APIs or executing functions. As LLM applications become more powerful and autonomous, their ability to correctly use APIs becomes critical. Similar to software libraries, APIs define the correct way to access desired functionality through structured calls. Inspired by this similarity, we propose extending ReadMe.LLM to support tool use by generating LLM-targeted documentation for APIs. Specifying important details like endpoints, parameters, or expected responses can be formatted and optimized for LLMs. Notably, the API-Bank benchmark demonstrates the importance of LLMs understanding API structure and semantics to improve tool use accuracy [53]. ReadMe.LLM can bridge the current performance gaps by providing LLM-friendly API documentation.
5 Conclusion and Future Work
We present ReadMe.LLM, novel LLM-oriented documentation that provides relevant context about a software library to assist code generation. We evaluated different combinations and structures of context and tested these across the current leading LLMs. We presented the optimal ReadMe.LLM structure, which has the highest average accuracy across different models, and increases correctness up to 100%.
As engineers continue to turn to LLMs when facing roadblocks, library developers must make their content easily available and understandable to LLMs. Failure to do so will not only hinder engineers by producing unreliable code but also disadvantage smaller libraries, perpetuating a cycle of underutilization and inefficiency. ReadMe.LLM becomes essential for bridging the gap between library documentation and Generative AI assistance.
Looking ahead, we remain committed to enhancing this framework by exploring new components and optimizing the structure of context delivery. We are planning on investigating how this framework extends to tasks other than code generation, such as question & answering and code debugging. Additionally, in this paper, we focus on LLM chatbots, but ReadMe.LLM can be extended to co-pilots as well. With the rise of vibe-coding and the adoption of products like Cursor [54], improving the code generation capabilities directly in the editor is important. A co-pilot could recognize the ReadMe.LLM file within an imported module’s directory and utilize it to generate more relevant and accurate code for its user.
We welcome contributions from the community to advance this initiative and shape the future of LLM-library interactions. Please explore our website, readmellm.github.io, and we encourage you to contribute to online discussions.
Acknowledgments
We thank the National Science Foundation and especially the SpectrumX (AST-2132700) community for its support. We also thank the UC Berkeley College of Engineering’s Fung Institute, as well as Dr. Josh Sanz for helpful conversations.
Appendix A Appendix: Comparison of results with older models
A.1 DigitalRF’s Case Study with Older Models
In our original experiments on the DigitalRF case study, we evaluated Claude 3.5 Sonnet across all eight context combinations and ReadMe.LLM, with web search enabled. When Claude 3.7 Sonnet was released, we repeated the same experiments.
A clear pattern emerges: while both versions benefit from richer context, the older Claude 3.5 model exhibited a larger performance jump when provided with our ReadMe.LLM versus human-oriented documentation. Specifically, Claude 3.5’s accuracy moved from 20% (no context or ReadMe.md) to 80% with ReadMe.LLM—a 60% point gain—whereas Claude 3.7 improved from 60% (no context or ReadMe.md) to 100% with ReadMe.LLM—a 40% point gain. This indicates that ReadMe.LLM had a greater impact in bridging the knowledge gap in the earlier version of the model. As models become more capable (e.g., Claude 3.7), the baseline performance on human-oriented docs rises, but ReadMe.LLM continues to provide a consistent boost.
When a new Sonar model was released built on Llama 3.3 70B, we also repeated our DigitalRF experiments to compare it with the older Sonar based on Llama 3.1 70B. Although we did not evaluate ReadMe.LLM on the older Sonar (Llama 3.1 70B), the eight standard context combinations reveal some clear trends. Both Sonar 3.1 and Sonar 3.3 achieved 100% success when provided only with examples, underscoring the power of real-world snippets. Most notably, Sonar 3.1 hallucinated dramatically under richer prompts—success falls to 40% with ReadMe.md + Examples and 20% with the full mixed context—whereas Sonar 3.3 remained far more robust (100% and 80%, respectively). In both versions, ReadMe.md alone yielded only a 20% success rate, again showing that human-oriented documentation offers minimal benefit.
A.2 Supervision’s Case Study with Older Models
When Claude was upgraded from 3.5 Sonnet to 3.7 Sonnet, we also reran Supervision’s Task 1 case study experiments. Claude 3.5 failed always with no context or ReadMe.md alone, yet jumped to a perfect 100% as soon as it saw examples, functions, or any combined context—including ReadMe.LLM. In contrast, Claude 3.7 already achieved 60% with no context and 20% with ReadMe.md, and likewise hits 100% once provided with ReadMe.LLM. Similar to the point we made for DigitalRF’s case study, the improvement from human-oriented docs to ReadMe.LLM is far more dramatic for Claude 3.5 than for Claude 3.7.
Appendix B Appendix: ReadMe.LLM Instructions
1.
Determine Functions to Include: Select the most popular or important functions and classes that represent the library’s core purpose and domain context.
2.
Extract Function Signatures: Clearly include function and method definitions with parameters and return types
3.
Find/Create Code Examples of Each Function: Pair signatures with illustrative code examples that demonstrate real-world usage of the functions
4.
Interleave Description of Function, Function Signature, and Example: Structure the content using sections with XML tags, combining a description, the function signatures, and the examples in a parsable layout. Below is the structure of the core building block of our ReadMe.LLM:
<context_1> <context_1_description> </context_1_description> <context_1_function> </context_1_function> <context_1_example> </context_1_example> </context_1>
5.
Add rules to the top of the ReadMe.LLM file. Below are the rules we added to our ReadMe.LLM files:
Rules:
Rule number 1: When you’re unsure about something, ask the user what information you need.
Rule number 2: Reuse SuperVision’s functions and code when applicable.
Rule number 3: Consider library dependencies when generating code solutions.
Appendix C Appendix: ReadMe.LLM Content
C.1 Supervision’s ReadMe.LLM
The following is the full ReadMe.LLM for the Supervision library:
C.2 DigitalRF’s ReadMe.LLM
The following is the full ReadMe.LLM for the DigitalRF library: